OpenStudy (anonymous):

What is the equation of the line passing through (–6, 1) and perpendicular to the line y = –3x + 1 in slope-intercept form? y = 3x + 4 y = 1/3x + 3 y = –3x – 6 y =1/3 x – 5

OpenStudy (lopezking1):

y = -3x + 1 ---> the slope for this line is -3 Because the slope is -3, the slope of the perpendicular line would be the inverse reciprocal, which is 1/3 You want the line to pass through (-6, 1). The equation for the line is y = 1/3x + b Find b. Plug in -6 for x and 1 for y. 1 = 1/3(-6) + b 1 = -2 + b b = 3 So the equation for the perpendicular line is y = 1/3x + 3 B

OpenStudy (anonymous):

-3*1/3 should always equal -1 to find the slope to a perpendicular slope

OpenStudy (anonymous):

so the answer is in fact -3